Week 43: October 19 - 26 Last updated:   19-Oct-20 12:10
Maintenance this week Date Time & Length Production resumes
520 MeV Tuesday, Oct. 20 starting at 08:30 for 31 hr, followed by 2 hours of start up Oct. 21 @ 17:30
ISAC Tuesday, Oct. 20 Oct. 21 @ 17:30
E-LINAC Tuesday, Oct. 20 E-Hall will be open on Tuesday, closed the rest of the week. -
                   
GROUP ACTIVITY Fault Report Work Permit Completed
ATG STF water package work in SAB3 520 MeV - C2020-10-20-8 y
STF water package flow measurements C2020-10-21-9 y
1AQ15 flush 520 MeV -
Beamlines Block moves at M9 & check polarities of M9Q3-7 520 MeV - C2020-10-20-5, C2020-10-20-9 y
1VB1 flush,  1V water manifolds maintenance (TBD) 520 MeV
Beam Development ISIS studies [Wednesday] 520 MeV - C2020-10-21-8 y
Computing and Networking
Controls Check water system TCs 520 MeV 13214
Investigate TG155 & 165 timing out 520 MeV 13487 C2020-10-20-11 n - no time
Investigate I1 GND side interlocks flooding msg reader 520 MeV 13453, 13450
Test SSH server restart on DEVBN2 520 MeV -
Investigate noisy LE1 signals 520 MeV 13490
Cryogenics and Vacuum Cyclotron cryopanel defrost 520 MeV - C2020-10-20-1 y
Prepare E-LINAC 2 K for diesel generator tests and recover E-LINAC - y
Pump down 1U 520 MeV - - y
Replace 2C turbo 520 MeV - C2020-10-21-11 y
Diagnostics Water Cooled Probe water line flush 520 MeV 13387 C2020-10-21-5 y
Load more combo foils for Ex1 520 MeV - C2020-10-21-3 y
Take pictures of South Periscope in prep for SD 2020-2021 replacement 520 MeV - C2020-10-21-6 y
Experimental Support M20 Cryopump operation observation.  520 MeV C2020-10-20-10 y
ISAC Target Hall ITE -> ITW (TBD) ISAC - - -
ISIS I1 backplate change [Tuesday] 520 MeV - C2020-10-20-2 y
Cryo014 work 520 MeV 13497 C2020-10-21-7 y
Magnets and Power Supplies Check MMPS control cabinet [MM off, Tuesday 12:00-15:00] 520 MeV - - n
Repair 1AQ9 polarity switcher 520 MeV
Repair MM power supply 520 MeV - C2020-10-20-7 y
Operations Sump checks All - - y
SCBCD & SCRM Tests 520 MeV - - y
2A3/ITE -> 2A2/ITW (TBD) 520 MeV/ISAC - - not needed
Installation of chamber and HV power supply at IIS ISAC - -
Turn on and test 1U devices, safety critical device testing 520 MeV - - n
Remote Handling
RF
Safety Tour vault and tunnels 520 MeV - C2020-10-21-10 y
Site Services SAB3 pump check 520 MeV - - n
Mechanical, Civil/Structural, Electrical, FAS Diesel generator load switch tests All - C2020-10-20-3 y
ISAC Chiller-1 preventative maintenance ISAC - y
Cu-ALCW sidestream dissolved oxygen sensor replacement (continued) 520 MeV 13448 - not done
Change cyclotron vault exhaust filters 520 MeV 13471 C2020-10-20-6 y
Change Cu-ALCW deionizer loop resin and filters 520 MeV 13481, 13473 C2020-10-21-1 y
Bring new cooling towers online + final commissioning 520 MeV - C2020-10-20-4 y - still some tests to do
E-LINAC HAW system work E-LINAC/520 MeV - y - to be continued
Switch RF HVAC from A/C backup to ME-141 520 MeV - C2020-10-20-12 y
Repair/replace ground water check valve 520 MeV C2020-10-22-2 n - old valve removed, new on order